# Can Masked Autoencoders Also Listen to Birds?

### v3

@@ 7 Conclusion and Future Work
Conclusion. In this work, we addressed the limitations of general-purpose SSL models in audio classification. We demonstrated the efficacy of a holistically adapted, domain-specific masked image modeling pipeline for bird sound classification: We revised the entire training process, including pretraining (e.g., replacing AudioSet with BirdSet), fine-tuning (e.g., adding prototypical pooling), frozen representations (e.g., utilizing prototypical probing), leading to the development of our Bird-MAE model. Bird-MAE achieves novel state-of-the-art performance on the BirdSet multi-label classification benchmark, strongly outperforming the general-purpose Audio-MAE and prior best-performing supervised models. Our findings highlight that while domain-specific pretraining is crucial, the full benefits of such adaptations become particularly evident when leveraging frozen representations. Specifically, our parameter-efficient prototypical probing substantially narrows the gap to full fine-tuning to 3%${}_{\text{p}}$ MAP on average across downstream tasks. It boosts it up to 37%${}_{\text{p}}$ compared to linear probing. These results underscore the importance of domain-aware pretrained features and effective probing methods. Furthermore, Bird-MAE with prototypical probing delivers strong few-shot performance, offering an efficient alternative for resource-constrained bioacoustic applications. Our study shows that achieving optimal results in more fine-grained audio tasks such as bioacoustics requires moving beyond generic SSL approaches towards holistic, domain-aware pipeline adaptations.
Future work and limitations. Several promising research directions emerge from our findings and the limitations of this work. First, exploring the generalizability and broader applicability of prototypical pooling and probing beyond bird sound classification is crucial. Learning discriminative prototypes and leveraging spatial feature maps is not intrinsically tied to bioacoustics. It suggests the potential for enhancing downstream task adaptation across SSL methods and modalities, including general audio, vision, and text. For instance, our preliminary results in Table 8 indicate that our prototypical pooling also notably improves the fine-tuning performance of the base Audio-MAE on AS-20k. Notably, this simple pooling replacement not only surpasses the original Audio-MAE performance reported by Huang et al. (2022)
Model AS-20k
A-MAEbase 37.3
A-MAEproto 38.6
BEATsiter2 38.3
Table 8: Impact of prototypical pooling on fine-tuning in general audio.
but also outperforms the advanced BEATs model (Chen et al., 2023) on this task. This result hints at the potential for prototypical layers to unlock greater performance from existing SSL backbones across different audio domains. Second, the learned prototypes (see Figure 5) can be further exploited beyond performance improvements: analyzing prototypes could improve model robustness, add model-based event detection, and provide insights for domain experts. Lastly, the strong few-shot performance of Bird-MAE with prototypical probing highlights its potential for real-world applications, particularly in scenarios where acquiring large amounts of labeled data is prohibitive or computing resources are limited. Our work contributes to developing more effective domain-specific SSL in bioacoustics and other challenging low-resource domains.
@@ A.1 Pretraining Data
[…]
@@ A.3 Full Downstream Data
The BirdSet dataset provides file-level (weak) labels for recordings that may contain multiple vocalization events, necessitating downstream processing to generate task-specific datasets (Rauch et al., 2024). Our preliminary experiments on small focal validation sets on all downstream tasks and models indicated that a uniform sampling strategy across all BirdSet’s tasks is suboptimal. Consequently, we adopted tailored approaches: For the datasets HSN, UHH, and NBP that have the lowest class counts across tasks, we utilized BirdSet’s inherent sampling strategy by adding a species cap of 500 and extracting a maximum of 5 events per recording, following Rauch et al. (2024). More details are available in the BirdSet implementations. Conversely, the datasets SNE, POW, NES, PER, and SSW are characterized by larger data volumes and higher class counts. We found it more advantageous to employ our few-shot sampling approach with $k=64$ to further reduce label noise and class imbalance. This proved beneficial for experiments involving both fine-tuning and frozen representations across all models. The final number of samples curated for each dataset in our study is detailed in the main text in Table 2.
@@ Appendix B Metrics
This appendix provides a detailed description of the evaluation metrics employed to assess model performance throughout this study. We evaluate model performance in the main paper with the mean average precision (MAP) as the metric in multi-label classification (Huang et al., 2022; Chen et al., 2023; 2024a; Rauch et al., 2024). In the additional results in Appendix E, we also report the area under the receiver operating characteristic curve (AUROC), and top-1 accuracy (T1-Acc), following the multi-label benchmark from BirdSet.
  • MAP, also referred to as class-wise MAP (cmAP) (Rauch et al., 2024), first calculates the average precision (AP) for each class $c$ independently and then computes the macro-average of these AP scores across all $C$ classes:
@@
  • T1-Acc assesses whether the class assigned the highest predicted confidence score by the model is among the set of true labels for a given instance (Rauch et al., 2024):
[…]
@@ D.1 Augmentations
Our data augmentation pipeline, applied during to all experiments, including few-shot multi-label classification and probing variants, is adapted from the strategies outlined in BirdSet (Rauch et al., 2024). We empirically tuned the application probability for each selected augmentation based on preliminary experiments on the validation data HSN. A key component of our pipeline is waveform-level mixup, implemented using TorchAudiomentations (Jordal et al., 2024), which demonstrated superior performance compared to spectrogram-based or standard linear mixup. For augmentations requiring external audio, such as background noise addition and no-call mixing, we utilized environmental recordings sourced from BirdSet’s VOX dataset. Additional waveform and spectrogram-level augmentations, along with their specific parameters, are detailed in Table 9.
@@ D.2 Hyperparameters
Model Training (Fine-Tuning / Probing on full data)
@@ E.2 Few-Shot Multi-Label Classification
Table 12: Frozen representation results for prototypical, linear and attentive probing on our few-shot multi-label classification benchmark (MAP, AUROC, T1-Acc.). Comparison of our best-performing Bird-MAE-L model with few-shot training data, following the evaluation protocol of (Rauch et al., 2024). Best results are highlighted. This complements Table 7 from the main text.

### v4

@@ 7 Discussion and Limitations
Shortfall of general-purpose MAEs. While our study demonstrates that a domain-specific and holistically adapted MAE pipeline can excel, there could be other reasons for the performance gap. Our work leaves open questions regarding other factors, such as (1) the model scale or the (2) the SSL objective. Regarding model scale, one might hypothesize that simply using a vastly larger general-purpose model would suffice. However, evidence suggests this approach has limitations. We observe diminishing performance returns when scaling our own adapted model from a ViT-Large to a ViT-Huge backbone. This aligns with findings from benchmarks like HEAR (Turian et al., 2022), where even billion-parameter general audio models have been shown to underperform on specific bioacoustic tasks. Regarding the SSL objective, another hypothesis is that the MAE reconstruction task is inherently less suited for this fine-grained domain than other methods. Evidence suggests the challenge is broader than one specific method. Our study shows that a holistically adapted Bird-MAE notably outperforms other domain-adapted SSL models that use different objectives, including the contrastive SimCLR and the masking-based AVES Table 6. This aligns with findings from BIRB bioacoustic benchmarks (Hamer et al., 2023), where the YamNET model (weak clip-level training) also exhibits performance limitations. Collectively, this suggests that while model scale and SSL objective influence feature quality, a comprehensive, domain-specific adaptation of the entire pipeline appears to be a dominant factor for achieving SOTA performance in this fine-grained domain. Nevertheless, a systematic cross-paradigm comparison on a standardized benchmark like BirdSet remains a valuable direction for future work to definitively isolate the impact of different SSL objectives and model sizes.
Prototype interpretability. The interpretability of the learned prototypes by heat map visualizations provide a notable value beyond performance metrics. Heinrich et al. conduct a comprehensive analysis demonstrating that prototypes can learn to represent human-interpretable sound patterns, such as distinct call types within a single bird species (as we illustrate exemplary Figure 5). While a detailed interpretability analysis of our specific Bird-MAE representations is beyond the scope of this performance-focused study, it represents an important direction for future research. For instance, the ability to inspect what a model has learned offers opportunities for human-in-the-loop learning, where ecologists could validate or refine prototypes to correct model errors and potentially further enhance performance in data-scarce scenarios (Rauch et al., 2024a). This synergy between strong few-shot performance and interpretability could enable the development of robust and transparent monitoring of of rare species with minimal labels.
8 Conclusion
@@
In this work, we addressed the limitations of general-purpose SSL models in audio classification. We demonstrated the efficacy of a holistically adapted, domain-specific masked image modeling pipeline for bird sound classification: We revised the entire training process, including pretraining (e.g., replacing AudioSet with BirdSet), fine-tuning (e.g., adding prototypical pooling), frozen representations (e.g., utilizing prototypical probing), leading to the development of our Bird-MAE model. Bird-MAE achieves novel state-of-the-art performance on the BirdSet multi-label classification benchmark, strongly outperforming the general-purpose Audio-MAE and prior best-performing supervised models. Our findings highlight that while domain-specific pretraining is crucial, the full benefits of such adaptations become particularly evident when leveraging frozen representations. Specifically, our parameter-efficient prototypical probing substantially narrows the gap to full fine-tuning to 3 pp mAP on average across downstream tasks and boosts it up to 37 pp over linear probing. These results underscore the importance of domain-aware pretrained features and effective probing methods. Furthermore, Bird-MAE with prototypical probing delivers strong few-shot performance, offering an efficient alternative for resource-constrained bioacoustic applications. Our study shows that achieving optimal results in more fine-grained audio tasks such as bioacoustics requires moving beyond generic SSL approaches towards holistic, domain-aware pipeline adaptations.
@@ A.1 Pretraining Data
We derive our pretraining set from BirdSet’s XCL collection using the provided file-level species labels, the event detector and the sampling algorithm from BirdSet (Rauch et al., 2024b). Since each recording may contain multiple bird-call events, we first split every recording into its detected events. To mitigate class imbalance, where some species have many events and others few, we enforce three constraints: (a) a maximum number of events per species, (b) a per-recording event cap, and (c) at least one event per recording. We implement these rules via a simple sampling algorithm that iteratively trims over-represented recordings until all class and event limits are met. We set the number of maximum events per species to 500 and a per-recording event cap to 2. This leads to our XCL-1.7M pretraining dataset. The uncurated dataset $\text{{XCL-3.4M}}_{\text{{R}}}$ contains all the detected events.
@@ A.3 Full Downstream Data
The BirdSet dataset provides file-level (weak) labels for recordings that may contain multiple vocalization events, necessitating downstream processing to generate task-specific datasets (Rauch et al., 2024b). Our preliminary experiments on small focal validation sets on all downstream tasks and models indicated that a uniform sampling strategy across all BirdSet’s tasks is suboptimal. Consequently, we adopted tailored approaches: For the datasets HSN, UHH, and NBP that have the lowest class counts across tasks, we utilized BirdSet’s inherent sampling strategy by adding a species cap of 500 and extracting a maximum of 5 events per recording, following Rauch et al. (2024b). More details are available in the BirdSet implementations. Conversely, the datasets SNE, POW, NES, PER, and SSW are characterized by larger data volumes and higher class counts. We found it more advantageous to employ our few-shot sampling approach with $k=64$ to further reduce label noise and class imbalance. This proved beneficial for experiments involving both fine-tuning and frozen representations across all models. The final number of samples curated for each dataset in our study is detailed in the main text in Table 2.
@@ Appendix B Metrics
This appendix provides a detailed description of the evaluation metrics employed to assess model performance throughout this study. We evaluate model performance in the main paper with the mean average precision (MAP) as the metric in multi-label classification (Huang et al., 2022; Chen et al., 2023; 2024a; Rauch et al., 2024b). In the additional results in Appendix E, we also report the area under the receiver operating characteristic curve (AUROC), and top-1 accuracy (T1-Acc), following the multi-label benchmark from BirdSet.
  • MAP, also referred to as class-wise MAP (cmAP) (Rauch et al., 2024b), first calculates the average precision (AP) for each class $c$ independently and then computes the macro-average of these AP scores across all $C$ classes:
@@
  • T1-Acc assesses whether the class assigned the highest predicted confidence score by the model is among the set of true labels for a given instance (Rauch et al., 2024b):
    $\text{T1-Acc}=\frac{1}{N}\sum_{n=1}^{N}\mathbb{I}\{\hat{y}^{(\text{top})}_{n}\in Y_{n,:}\},$
@@ D.1 Augmentations
Our data augmentation pipeline, applied during to all experiments, including few-shot multi-label classification and probing variants, is adapted from the strategies outlined in BirdSet (Rauch et al., 2024b). We empirically tuned the application probability for each selected augmentation based on preliminary experiments on the validation data HSN. The spectrogram time-frequency masking is similar to SpecAugment (Park et al., 2019). However, we omit the time-warping component of SpecAugment, which did not improve validation performance in our setting. A key component of our pipeline is waveform-level mixup, implemented using TorchAudiomentations (Jordal et al., 2024), which demonstrated superior performance compared to spectrogram-based or standard linear mixup. For augmentations requiring external audio, such as background noise addition and no-call mixing, we utilized environmental recordings sourced from BirdSet’s VOX dataset. Additional waveform and spectrogram-level augmentations, along with their specific parameters, are detailed in Table 9.
@@ D.2 Hyperparameters
[…]
@@ E.2 Few-Shot Multi-Label Classification
Table 12: Frozen representation results for prototypical, linear and attentive probing on our few-shot multi-label classification benchmark (MAP, AUROC, T1-Acc.). Comparison of our best-performing Bird-MAE-L model with few-shot training data, following the evaluation protocol of (Rauch et al., 2024b). Best results are highlighted. This complements Table 7 from the main text.
@@ Appendix F Additional Ablations
Appendix G Generalizability Study on MeerKAT
To investigate the generalizability of our findings beyond avian bioacoustics, we conduct a preliminary study on the MeerKAT dataset (Schäfer-Zimmermann et al., 2024), which contains multi-label meerkat vocalizations. This small-scale study was designed to test two hypotheses: (a) that our domain-specific Bird-MAE provides transferable benefits to other fine-grained bioacoustic tasks, even when used as a frozen feature extractor, and (b) that prototypical probing remains superior to linear probing in this new domain.
Dataset and preprocessing. The MeerKAT dataset consists of 10-second multi-label audio clips with eleven unique call types originally sampled at 8kHz. As the dataset does not provide an official train-test split, we follow the protocol from Schäfer-Zimmermann et al. (2024) by randomly sampling 20% of the data to create a test dataset. To ensure efficient experimentation while demonstrating relative performance gains, we randomly sample 50% of the remaining data for our training set with a small validation split. During data loading, all audio recordings were upsampled to to match the pretraining specifications of our models. For compatibility with Bird-MAE’s 5-second input window, the 10-second clips from the MeerKAT dataset were segmented into two non-overlapping 5-second chunks.
Training and augmentation. For this study, we evaluate frozen representations from both the general-purpose Audio-MAE and our domain-specific Bird-MAE. We apply linear probing and our proposed prototypical probing to both models. The hyperparameters for the probing heads (e.g., learning rate, number of prototypes) are kept consistent with those used for the main experiments. We use a minimal augmentation strategy, applying only multi-label mixup with a probability of 0.5 during training.